OpenStudy (dr.gray):

Which number sentence is true? {A}. 0.44 > 0.44 {B}. -5.55 > 5.55 {C}. -4.5 > -3.60 (D). -3 5/8 > -4 3/5 (Will Fan & Medal) {}= false ()= true

Awolflover1 (awolflover1):

alright 0.44 > 0.44 is false

Awolflover1 (awolflover1):

-5.55 > 5.55 is false too its the same thing with 5.55

Awolflover1 (awolflover1):

now we're down with 2 which one is true?

Awolflover1 (awolflover1):

\[−4.5≯−3.6\]

OpenStudy (dr.gray):

C

Awolflover1 (awolflover1):

C is false

OpenStudy (dr.gray):

Darn

Awolflover1 (awolflover1):

\[−4.375>−8.6\] :)

Awolflover1 (awolflover1):

hmm ..

OpenStudy (dr.gray):

this is the next question

Awolflover1 (awolflover1):

this is a hard one. :\ make another post Dr. ;-;

OpenStudy (dr.gray):

how about this one

OpenStudy (mathstudent55):

What is \(10^2\) equal to?

OpenStudy (dr.gray):

100

OpenStudy (mathstudent55):

Good. Then \(5.634 \times 10^2 = 5.634 \times 100\) What is that equal to?

OpenStudy (dr.gray):

563.4

OpenStudy (mathstudent55):

Good. Now compare with the options.

OpenStudy (mathstudent55):

You have the number 563.4 Look at the options. a. Is 563.4 between 0.0562 and 0.0563? b. Is 563.4 between 0.0563 and 0.0564? 563.4 is more than 500. 0.0562, 0.0563, and 0.0564 are a little more than half. 563.4 is way, way more than those numbers. a. and b. are out

OpenStudy (mathstudent55):

a. Is 563.4 between 562 and 563? b. Is 563.4 between 563 and 564? How does 563.4 compare with 562, 563, and 564?

OpenStudy (mathstudent55):

In the middle problem you posted above, you have to write three numbers in order, from least to greatest. Take the absolute value of the negative number. You will end up with a number that is in the units, between 0 and 10. One number in the tens, between 10 and 100. The third number is in the hundreds, between 100 and 1000. If you think that way, it will be easy to write them in order.